OpenStudy (idani):

what is the solution of the equation? eliminate any extraneous solutions. (-2x+6)^1/5 =-8+10x)^1/5

OpenStudy (lilah):

i got you hold up

OpenStudy (idani):

A. 2/3 B. 6/7 C. -1/4 D. 7/6

OpenStudy (lilah):

damn sorry i dont know i used math way and it says there's no solutuions

OpenStudy (lilah):

use math way its some calculator online and it solves everything u could think of

OpenStudy (idani):

Okay thank you ☺️

TheSmartOne (thesmartone):

Is that \(\Large (-2x+6)^{\frac{1}{5}} =(-8+10x)^{\frac{1}{5}}\) If so, use this: IF \(\Large a^c = b^c\) THEN \(\Large a = b\)

OpenStudy (idani):

Thank you☺️

TheSmartOne (thesmartone):

Anytime! :)